40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 18 20:17:17 2014, Jenne, Update, LSC, ALS noisy ALS_1kHzRMSnoise.pdfALS_oolNoise.pdf
    Reply  Mon Apr 21 22:53:16 2014, manasa, Update, LSC, ALS noise ALSX_FreeInLoop.jpgALSXY_inLoop.jpg
       Reply  Tue Apr 22 01:11:42 2014, Jenne, Update, LSC, TRY 60Hz noise TRY_60Hz_noise_21Apr2014.pdf
          Reply  Tue Apr 22 21:54:50 2014, manasa, Update, LSC, TRY 60Hz noise 60HzTRY.jpg
             Reply  Tue Apr 22 22:49:10 2014, rana, Update, LSC, TRY 60Hz noise 
                Reply  Wed Apr 23 19:58:00 2014, manasa, Update, LSC, TRY 60Hz noise 
                   Reply  Tue May 6 18:32:14 2014, steve, Update, LSC, TRY 60Hz noise hunt ETMY-ISCT_EISOL.jpg
                      Reply  Wed May 7 14:01:36 2014, steve, Update, LSC, TRY 60Hz noise hunt 
                         Reply  Thu May 8 17:25:17 2014, steve, Update, LSC, TRY 60Hz noise hunt 
                         Reply  Fri May 16 16:08:12 2014, steve, Update, LSC, Isolating base plates ISObaseplates.jpg
                Reply  Wed Apr 23 23:48:30 2014, manasa, Update, LSC, Y end whitening board IMG_1378.JPGIMG_1379.JPGIMG_1378.JPG
                   Reply  Thu Apr 24 14:23:09 2014, not manasa, Update, LSC, Y end whitening board 
                      Reply  Thu Apr 24 23:55:31 2014, Koji, Update, LSC, Y end whitening board P4245550.JPG
                         Reply  Fri Apr 25 10:43:57 2014, Koji, Update, LSC, (Fixed) Y end whitening board D990399-B_40m.pdfP4245552.JPGP4245553.JPGP4245551.JPG
                            Reply  Thu Dec 11 18:12:46 2014, ericq, Update, LSC, (Fixed) Y end whitening board 
                               Reply  Thu Dec 11 22:44:23 2014, rana, Update, LSC, QPD screens 
          Reply  Thu Apr 24 00:11:35 2014, Jenne, Update, LSC, Yend shutter back. 
       Reply  Sun Apr 27 20:26:19 2014, Koji, Update, LSC, Phase Tracker servo characterization ALSX_PTTF.pdfALSY_PTTF.pdf
          Reply  Sun Apr 27 21:30:59 2014, Koji, Update, LSC, ALS servo characterization ALSX_OLTF.pdfALSY_OLTF.pdf
             Reply  Mon Apr 28 10:24:10 2014, Koji, Update, LSC, error signal characterization ALSX_SPE.pdfALSY_SPE.pdf
                Reply  Mon Apr 28 10:34:51 2014, Koji, Update, LSC, new ALS servo design ALSX_OLTF_new2.pdfALSY_OLTF_new2.pdf
                   Reply  Mon Apr 28 10:48:48 2014, Koji, Update, LSC, new ALS servo design: comparison ALSX_OLTF_new.pdfALSY_OLTF_new.pdf
                      Reply  Mon Apr 28 10:59:54 2014, Koji, Update, LSC, New ALS servo design: expected error signals ALSX_SPE_new.pdfALSY_SPE_new.pdf
                         Reply  Mon Apr 28 11:08:11 2014, Koji, Update, LSC, New ALS servo design: expected error signals 140421_ALS_servo.zip
Message ID: 9854     Entry time: Fri Apr 25 10:43:57 2014     In reply to: 9852     Reply to this: 10785
Author: Koji 
Type: Update 
Category: LSC 
Subject: (Fixed) Y end whitening board 

I went to WB and found the last spare module of D990399 revB. We need to thank Frank for his foresight.

The original (=broken) board had various modifications from this revB.
I had to check the schemaric diagram and the difference between the boards and migrate some of the SMD components from left to right.


Here is the deciphered features of the QPD whitening board:
- The input stage is a VGA amp (AD602). It has the internal input impedance of 100 Ohm. The series resister
  of 909 Ohm gives us 1/10 voltage division! It is more tricky as the QPD (D990272) has the output impedances of 50Ohm
  (for the both side of the differential out) and on resistance of MAX333A. So it could have been deviated by ~10% from the nominal.

- Variable gain control: The input has 1/10 voltage division. The gain is fixed at the unity. In total the gain of the variable control stage is 1/10.
  This gives us the gain range of +42dB/-22dB for +10V/-10V. The actual range is limited to be -10~30dB.

- Whitening stages. Each channel has two sets of the whitening path and the bypass path.
  They could be switched by binary control inputs but I permanently enabled the whitening by pulling the MAX333 control inputs to the ground.
  The whitening zero and pole are at 4.02Hz and 40.6Hz.

  Each bypass path has an additional cap of 220pF in parallel to 35.7kOhm (R101 and R103 for CH1), resulting in the pole at 20.2kHz.
  Each whitening paths had a 5.6nF cap (C53 and C64). This cap was replaced with 350pF, resulting in the move of the pole freq from 800Hz to 12.7kHz.

- There are two anti-aliasing stages which were designed for 2kHz sampling rate. They are identical sallen key 2nd-order LPFs with fc=766Hz and Q=0.74 (~ butterworth).
  As all of these caps were removed, they are just voltage followers now.

- The final stage (AD620) has the gain resister of 16.5k. The gain is 1+(49.4k/16.5k) = 3.99.

- The 4pin lemo connector (J8) was removed from the board. We instead installed an isolated BNC connector on the panel for the thorlabs PD serving as the high gain PD.

- There is a daughter board for the high gain PD. This seems to be the butterworth low pass filter with fc=~30kHz.
  The differential output of the daughter board is connected to pin 17 and 18 of J10 (S5 Out and Rtn).

- The input of the daughter board is differential (AD620). Therefore the LEMO connectros next to the BNC were wrapped with Kapton tapes for isolation.

Board test at the workbench.

- The test required two dual power supply as the unit requires +/-5V and +/-15V.

- The four channels were tested with the signal injection. 1kHz input yielded 20mVpp across the AD602 input. The output of the 1st whitening stage was
  60mVpp. This makes sense as the gain of the AD620 is -10dB (1/10 and 10dB). The output of the 2nd whitening stage was 600mVpp.
  Finally the output of the output stage was confirmed to be 2400mVpp. This was confirmed for four channels.

- The daughter board output was also checked. The gain is the unity and flat upto ~10kHz.

Board installation

- Jenne installed the module. This time there was no smoke.


Gain mystery

- It was not sure how the whitening gains have been given.

- The corresponding database entry was found in /cvs/cds/caltech/target/c1auxey/ETMYaux.db as

grecord(ao,"C1:ASC-QPDY_S1WhiteGain")
grecord(ao,"C1:ASC-QPDY_S2WhiteGain")
grecord(ao,"C1:ASC-QPDY_S3WhiteGain")
grecord(ao,"C1:ASC-QPDY_S4WhiteGain")

- The gains for S2-S4 were set to be 30. However, C1:ASC-QPDY_S1WhiteGain was set to be 8.62068.
And it was not writable.

- After some investigation, it was found that the database was wrong. The DAC channel was changed from S100 to S0.
The corrected entry is shown here.

grecord(ao,"C1:ASC-QPDY_S1WhiteGain")
{
        field(DESC,"Whitening gain for QPDY Seg 1")
        field(DTYP,"VMIVME-4116")
        field(OUT,"#C0 S0 @")
        field(PREC,"1")
        field(EGUF,"42")
        field(EGUL,"-22")
        field(EGU,"dB")
        field(LINR,"LINEAR")
        field(DRVH,"30")
        field(DRVL,"-10")
        field(HOPR,"30")
        field(LOPR,"-10")
}

- Once c1auxey was rebooted, the S1 whitening gain became writable. Now all of the channels were set to be +30dB (max).

 

Attachment 1: D990399-B_40m.pdf  343 kB  Uploaded Fri Apr 25 16:26:21 2014  | Hide | Hide all
D990399-B_40m.pdf D990399-B_40m.pdf D990399-B_40m.pdf
Attachment 2: P4245552.JPG  4.663 MB  Uploaded Fri Apr 25 18:50:08 2014  | Hide | Hide all
P4245552.JPG
Attachment 3: P4245553.JPG  4.695 MB  Uploaded Fri Apr 25 18:50:45 2014  | Hide | Hide all
P4245553.JPG
Attachment 4: P4245551.JPG  227 kB  Uploaded Fri Apr 25 18:58:57 2014  | Hide | Hide all
P4245551.JPG
ELOG V3.1.3-